

********************************************************************
UARS WEEKLY STATUS REPORT - 26 Aug 2005
********************************************************************
On 21 August 2005 UARS suffered a shorted cell in Battery 2.
This event was one of the predefined trigger points referenced
in the UARS End of Mission plan. Planning for the UARS
decommissioning has been started. UARS will be in forward
flight until 17 September. Once UARS is yawed, de-orbit
burns can commence. Detailed plans are being prepared.
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005099 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005099 has been processed and is waiting for DAAC archival.
HEPSB data files are version 1. HEPSB data files between 1991274 and
1991366 have been processed and are waiting for DAAC archival.
Processing of all PEM DAAC data files has been suspended until the
DAAC catches up with archiving all processed data files. PEM data from
1991272 through 2005169 are available from SwRI. Users of SDDAS have this
data currently available. Please update your meta data to reflect the
current PEM data holdings.
AXIS1 data file formats are beginning design. Preliminary AXIS1
data format being coded for test. If these tests become successful, the
AXIS1 file format will be coded and tested.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 13 August 2005 (Orbit 76204, GMT Day 225) through
19 August 2005 (Orbit 76309, GMT Day 231).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 226/1354 HALOE POWER ON
The beta angle ranged between the angles of 10.7 degrees to 38.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 66 degrees on 29 August 2005 (DOY 242).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 9.73 to 5.83 degrees C. Additional
battery performance data for each battery on 12 August 2005 and 19
August 2005 are provided in Appendix D.
The clock error ranged from 11.3 to -12.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 10
HALOE = 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Zero (0) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to closely monitor the batteries. The
differential voltages on battery 2 have stabilized from last week.
The max temperature on battery 2 has slowly started increasing again.
An attempt to lower the temperature was made on 13 August 2005 by
changing VT levels from LV6 to LV5. It was unsuccessful and the
charging level was promptly returned to LV6.
MEETINGS
The FOT was asked to attend an ERBS End Of Mission meeting on 19
August 2005 to answer a few questions that Flight Dynamics Facility
(FDF) had in regards to the way in which UARS was planning on
executing the decommissioning plans.
A UARS De-boost kickoff meeting is scheduled for 22 August 2005. This
meeting will attempt to get all the required support groups together
so that everyone can begin working on their appropriate tasks to
support the De-boost and eventual termination of UARS.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 September 2005 (DOY 260).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 August 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
76214 225/1511 05-257 MPS VT L6 TO L5
76216 225/1831 05-258 MPS VT L5 TO L6
76228 226/1354 05-259 HALOE POWER ON PT 1
76231 226/1906 05-260 HALOE POWER ON PT 2
76231 226/1906 05-261 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76250 228/0046 05-262 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
76253 228/0631 05-263 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
76267 229/0443 05-264 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
76269 229/0657 05-265 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76274 229/1456 05-266 HALOE RTS 19
76286 230/1117 05-267 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
76293 230/2114 05-268 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
76302 231/1242 05-269 MPS VT L6 TO L5
76306 231/1855 05-269 MPS BATSOC
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 August 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 August 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
76231 226/1919 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76250 228/0051 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
76253 228/0646 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
76267 229/0446 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
76269 229/0702 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76286 230/1120 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
76293 230/2120 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 August 2005
12 August 2005 - GMT Day 224 Beta = 10.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -3.05/ -4.31 +00.0/+00.0 04.64 0.00 0.0
2 -89.6/-280.0 +11.07/+10.06 +31.6/-16.4 26.24 1.09 20.9
44.8/-700.0 +4.18/+ 3.23 +20.8/-09.6 26.08 1.05 13.8
19 August 2005 - GMT Day 231 Beta = 38.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.42/ -3.68 +00.0/+00.0 04.6 0.00 0.0
2 -50.4/-347.2 +14.55/ +9.06 +31.6/-19.6 25.1 1.07 22.2
3 33.6/-700.0 +4.82/+ 3.23 +19.2/-12.0 25.0 1.04 13.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
76214 225/1511 LV6 -> LV5
76216 225/1831 LV5 -> LV6
76302 231/1855 LV6 -> LV5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 August 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-039 FOT 76162 222/0506 40451 WS Chain Failover
05-040 FOT 76176 223/0244 40465 TLM HITS - WEATHER AT GUAM
05-041 FOT 76183 223/1443 40470 TLM DROPOUT - WEATHER AT
GUAM
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.